9/17 A: Carbon has two naturally occurring isotopes: C-12 and C-13. C-12 has an abundance of 98.93% and C-13 has an abundance of 1.07%. Calculate the average atomic mass of carbon.


1
Expert's answer
2021-03-22T06:26:51-0400

To find the average atomic mass, you take the percentage of abundance of isotopes, multiply by the total mass of each isotope, and then together.


"(\dfrac{98.93}{100} × 12 ) + (\dfrac{1.073}{100} × 13) ="


"11.8716 + 0.13949 ="


"12.01109\ amu".